Integrate ln(2x+2) from -1<=x<=1


1
Expert's answer
2021-08-02T17:39:38-0400

"\int"-1 1 ln(2x+2) dx


Let 2x + 2 = 2t ..................Equation(1)


Differentiating both the sides we have


dx = dt and x varies from -1 "\rightarrow" 1 so by using Equation(1) t varies from 0 "\rightarrow" 2



Now the integral takes the form


"\int"0 2 ln(2t) dt


By using integration by parts and considering u = ln(2t) and v = 1, we have


[ ln(2t) t - "\int" "\dfrac{2}{2t}" . t dt ]0 2


[ ln(2t) t - "\int" 1 dt ]0 2


[ ln(2t) t - t ]0 2


On substituting the limits


2[ ln (4 )- 1]